OpenStudy (anonymous):

PLEASE HELP how do you convert X+2y<2 into slope intercept form

OpenStudy (anonymous):

ok so the slope intercept form is y=mx+b where m is the slope and b is the y intercept. so you want to move everything but the y over to the left side to begin with

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so once we do that it looks like this \[2y<2-x\]

OpenStudy (anonymous):

we then divide the whole thing by 2 to get y by itself

OpenStudy (anonymous):

so then we have \[y<1-\frac{ x }{ 2 }\]

OpenStudy (anonymous):

now all you have to do is switch the x and the b so it turns into \[y<-\frac{ 1 }{ 2 }+1\]

OpenStudy (anonymous):

and thats your final answer, does that make sense?

OpenStudy (anonymous):

Yes! thank you

OpenStudy (anonymous):

yep no problem!